Circular No. 6176

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995Q IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of another supernova (mag ∼ 18) found on an R plate taken by M. J. Drinkwater with the U.K. Schmidt Telescope on May 31.7 UT and located at α = 20^{h}16^{m}29^{s}.51, δ = −43^{o}24′24′′.6, which is 20′′.5 east and 5′′.1 south of the galaxy’s center (offsets measured from a CCD image obtained with the 1.0-m reflector at Siding Spring on June 2 by McNaught). Nothing appears at the position of the suspect either on the ESO R survey or the SERC J survey. Stars of mag 16.5 and 18 have the following respective position end figures: 27^{s}.57, 41′′.9; 29^{s}.34, 25′01′′.8. The latter star is ∼ 0.1 mag fainter than the supernova. R. A. Stathakis and S. Lee, Anglo-Australian Observatory, obtained a spectrogram (range 300–1020 nm) with the Anglo-Australian Telescope on June 1; the spectrum shows this to be a normal type-II supernova, ∼ 2 months old.

BL LACERTAE R. C. Vermeulen, California Institute of Technology (CIT); I. W. A.

Browne, Jodrell Bank and CIT; M. H. Cohen, CIT; R. W. Goodrich, Space Telescope Science Institute; and P. M. Ogle, A. C. S. Readhead, and H. D. Tran, CIT, write: “We report the clear detection of broad Hα line emission in a spectrum of BL Lac taken on May 21.49 UT at the Hale telescope; this emission was confirmed on June 1.56 at the Keck telescope. Preliminary reduction of the May 21 data shows Hα with EW = 0.75 nm and FWHM = 4000 km/s. [N II] at 658.4 nm is superimposed; it is unresolved (FWHM ≲ 600 km/s), has redshift z = 0.068 and EW = 0.1 nm. Hints of Hβ and [O III] 500.7-nm emission need verification and further analysis. Na I D is the most prominent absorption feature; other stellar absorption features may be present also. Neither spectrogram was taken in photometric conditions; crude flux calibration, as well as visual inspection of the guider field at Keck, suggest that the continuum is probably in a faint state.”

COMET 58P/JACKSON-NEUJMIN Unfiltered CCD total magnitude reported by J. V. Scotti, Kitt Peak

(0.9-m Spacewatch telescope): May 22.42 UT, 21.3. The comet appeared very nearly stellar, the image being “softer” than nearby field stars.

1995 June 2 Daniel W. E. Green